# Electromagnetic Inverse Scattering from a Single Transmitter

### v5

@@ C.2 Metrics
Following previous work (Liu et al., 2022), we evaluate the quantitative performance of our method using PSNR (Wang and Bovik, 2009), SSIM (Wang et al., 2004), Relative Root-Mean-Square Error (MSE) (Song et al., 2021) and IoU. For PSNR, SSIM and IoU, a higher value indicates better performance. For MSE, a lower value indicates better performance. MSE is a metric widely used in EISP, which is defined as follows:
@@ Appendix D Additional Experimental Results
This section provides additional experimental results, including statistical validation (Section˜D.1) and extended qualitative comparisons (Section˜D.2). The latter includes more visual comparisons with baseline methods, as well as supplementary ablation studies on noise robustness and training data size.
@@ D.2 Additional Qualitative Results
We present more qualitative comparison on Circular dataset (Luo et al., 2024) and MNIST dataset (Deng, 2012) under settings with different transmitter numbers $N$ and noise levels, as shown in Fig.˜11 to Fig.˜16. We also present qualitative comparison on IF (Geffrin et al., 2005) dataset under the single-transmitter setting, as shown in Fig.˜17. Our method achieves comparable or superior performance to SOTA methods in most cases under multiple-transmitter settings, as Fig.˜11, Fig.˜12, Fig.˜15, and Fig.˜16 indicate. As shown in Fig.˜12 and Fig.˜16, PGAN (Song et al., 2021) introduces noticeable background artifacts due to the lack of consideration of physics. And Fig.˜11 shows that Img-Interiors occasionally fails to converge due to local optima.
For single-transmitter setting ($N=1$), our method significantly outperforms all previous approaches across all datasets, as shown in Fig.˜13 and Fig.˜14. Conventional methods such as BP(Belkebir et al., 2005), Gs SOM (Chen, 2009), and 2-fold SOM (Chen, 2009) produce only blurry reconstructions. Deep learning-based methods such as BPS (Chen, 2018), Physics-Net (Liu et al., 2022), and PGAN (Song et al., 2021) tend to “hallucinate” the digit, leading to wrong reconstruction on the MNIST dataset. Img-Interiors (Luo et al., 2024) produces results with structural errors that deviate significantly from the true morphology. In contrast, our method can still produce reasonably accurate reconstructions of the relative permittivity under such a challenging condition. This enables practical applications with fewer transmitters, significantly reducing deployment costs while preserving reconstruction quality.
@@
We present more qualitative comparison on 3D MNIST (de la Iglesia Castro, ) dataset under the single-transmitter setting, as shown in Fig.˜18. Our method successfully approximates permittivity reconstruction even in this difficult setting, whereas Img-Interiors (Luo et al., 2024) fails. Similar to the 2D scenario, Img-Interiors suffers from degeneration and can only produce fixed patterns, failing to capture the shape of scatterers.
Additionally, we evaluate our method on the more complex 3D ShapeNet (Wu et al., 2015) dataset to demonstrate its generalization capability. As shown in Fig.˜19, our approach can successfully reconstruct various complex structures including airplanes, cars and tables under the same single-transmitter configuration, maintaining accurate permittivity distribution recovery. This demonstrates the strong potential of our method for practical real-world applications.
@@
We present qualitative results across multiple noise levels (ranging from 5% to 30%) on the MNIST (Deng, 2012) dataset, providing visual support for the noise robustness analysis. As shown in Fig.˜20, our method maintains high visual fidelity at low noise levels. More importantly, the method maintains stable reconstructions and preserves the essential structure of the scatterer across the entire tested noise range (5% to 30%), demonstrating a smooth and gradual degradation in quality that is fully consistent with the quantitative results.
@@
We present qualitative results across different training data scales (ranging from 25% to 100%) on the MNIST (Deng, 2012) dataset, offering visual insights into the impact of data volume on reconstruction performance. As shown in Fig.˜21 and Fig.˜22 , our model maintains satisfactory reconstruction integrity even with limited training data.
@@
Figure 17: Qualitative comparison under single-transmitter setting on IF dataset. The results are obtained with $N=1$ transmitter, without noise level. Colors represent the values of the relative permittivity.

### v6

@@ C.2 Metrics
Following previous work [26], we evaluate the quantitative performance of our method using PSNR [44], SSIM [43], Relative Root-Mean-Square Error (MSE) [39] and IoU. For PSNR, SSIM and IoU, a higher value indicates better performance. For MSE, a lower value indicates better performance. MSE is a metric widely used in EISP, which is defined as follows:
@@ Appendix D Additional Experimental Results
This section provides additional experimental results, including statistical validation (Sec.˜D.1), extended qualitative comparisons (Sec.˜D.2) and additional ablation studies on loss functions (Sec.˜D.3). The qualitative comparisons include more visual comparisons with baseline methods, while the ablations further analyze noise robustness, training data size, and the impact of different training losses.
@@ D.2 Additional Qualitative Results
We present more qualitative comparison on Circular dataset [27] and MNIST dataset [12] under settings with different transmitter numbers $N$ and noise levels, as shown in Fig.˜12 to Fig.˜17. Our method achieves comparable or superior performance to SOTA methods in most cases under multiple-transmitter settings, as Fig.˜12, Fig.˜13, Fig.˜16, and Fig.˜17 indicate. As shown in Fig.˜13 and Fig.˜17, PGAN [39] introduces noticeable background artifacts due to the lack of consideration of physics. And Fig.˜12 shows that Img-Interiors occasionally fails to converge due to local optima.
For single-transmitter setting ($N=1$), our method significantly outperforms all previous approaches across all datasets, as shown in Fig.˜14 and Fig.˜15. Conventional methods such as BP[4], Gs SOM [6], and 2-fold SOM [6] produce only blurry reconstructions. Deep learning-based methods such as BPS [7], Physics-Net [26], and PGAN [39] tend to “hallucinate” the digit, leading to wrong reconstruction on the MNIST dataset. Img-Interiors [27] produces results with structural errors that deviate significantly from the true morphology. In contrast, our method can still produce reasonably accurate reconstructions of the relative permittivity under such a challenging condition. This enables practical applications with fewer transmitters, significantly reducing deployment costs while preserving reconstruction quality.
@@
We present more qualitative comparison on 3D MNIST [11] dataset under the single-transmitter setting, as shown in Fig.˜18. Our method successfully approximates permittivity reconstruction even in this difficult setting, whereas Img-Interiors [27] fails. Similar to the 2D scenario, Img-Interiors suffers from degeneration and can only produce fixed patterns, failing to capture the shape of scatterers.
Additionally, we evaluate our method on the more complex 3D ShapeNet [46] dataset to demonstrate its generalization capability. As shown in Fig.˜19, our approach can successfully reconstruct various complex structures including airplanes, cars and tables under the same single-transmitter configuration, maintaining accurate permittivity distribution recovery. This demonstrates the strong potential of our method for practical real-world applications.
@@
We present qualitative results across multiple noise levels (ranging from 5% to 30%) on the MNIST [12] dataset, providing visual support for the noise robustness analysis. As shown in Fig.˜20, our method maintains high visual fidelity at low noise levels. More importantly, the method maintains stable reconstructions and preserves the essential structure of the scatterer across the entire tested noise range (5% to 30%), demonstrating a smooth and gradual degradation in quality that is fully consistent with the quantitative results.
@@
We present qualitative results across different training data scales (ranging from 25% to 100%) on the MNIST [12] dataset, offering visual insights into the impact of data volume on reconstruction performance. As shown in Fig.˜21 and Fig.˜22 , our model maintains satisfactory reconstruction integrity even with limited training data.
D.3 Additional Ablation Results
We experiment with TV loss [34] for smoothness and Perceptual loss [20] for structure in addition to MSE on single-transmitter setting. Tab.˜5 reports the quantitative results on the MNIST dataset under different noise levels. Overall, the auxiliary losses provide mixed improvements, while MSE alone already provides strong and stable performance, indicating that our model is relatively robust to the choice of loss function.
Table 5: Ablation of loss functions. We compare MSE with auxiliary TV loss [34] and perceptual loss [20] under the single-transmitter setting. Results are reported on the MNIST dataset under two noise levels: 5% and 30%.
Loss MNIST (5%) MNIST (30%)
MSE $\downarrow$ SSIM $\uparrow$ PSNR $\uparrow$ MSE $\downarrow$ SSIM $\uparrow$ PSNR $\uparrow$
Ours (MSE) 0.085 0.921 26.09 0.127 0.862 22.56
Ours (w/ TV) 0.084 0.917 26.00 0.124 0.859 22.74
Ours (w/ Percpt.) 0.084 0.923 26.20 0.130 0.861 22.49